# Experimental practical quantum tokens with transaction time advantage

### v1

@@ 5 Related work and extensions of our schemes
Like our implementation, users/clients in that of Ref. [35] rely on clock synchronization and position authentication, using networks which are potentially insecure. As noted above, these issues potentially affect the correctness of any quantum token scheme, including the present implementation of ours and that of Ref. [35]. If the client’s coordinates are incorrect, she may present a token in an incorrect region, i.e. one where no merchant or bank agent is present, or in a valid but unintended region. This means she may spend her token to acquire resources in the wrong space-time region, which may be to her disadvantage (for example, if she has a time- and location-dependent trading strategy). It also affects the privacy of any quantum token scheme. The client may present quantum money tokens in a region in the causal past of her intended region, which would allow the spoofer (e.g. the bank, or a third party) to learn her intended presentation region. Alternatively, she may present tokens in a region space-like separated from her intended region, which allows the spoofer exploitable information about her trading strategy in regions where she requires such information to be unavailable. Again, this may be disadvantageous when her trading strategy and those of adversaries (such as the spoofer) are time- and location-dependent and utilize all locally available information. We emphasize again that secure position authentication and time synchronization are solvable problems – assuming that some data, some locations and a reference clock can be trusted – although they are not experimentally addressed in our present implementation, nor in Ref. [35], nor indeed, as far as we are aware, in any implementation to date in the field of relativistic quantum cryptography (see e.g. [31, 32, 51, 52, 53]).
@@
Neither the present scheme nor that of Ref. [35] requires space-like separated presentation regions or other space-time constraints to ensure unforgeability. The reason we impose these constraints in the present implementation is to guarantee that the implementation demonstrates an advantage over classical alternatives. We emphasize again that in any form of quantum token scheme, unless the presentation and validation of quantum tokens can be carried out sufficiently fast to show an advantage in the transaction time over classical cross-checking schemes, there is no clear motivation to use any form of quantum tokens rather than this simple classical alternative.
We also emphasize that the security proof for unforgeability given in Ref. [14] considered arbitrary attacks that included arbitrary loss-dependent attacks, as does the new security analysis given in the present work. The analysis of Ref. [14] allowed an arbitrarily powerful dishonest user to receive all the quantum states transmitted by the bank and to report any subset of states as lost in the quantum channel. As discussed in Refs. [15, 14], freedom in reporting losses opens the potential loophole of multiphoton attacks, which compromises user privacy as a dishonest bank can use these attacks to obtain information about the user’s chosen measurement bases, and hence about the user’s chosen presentation regions. The implementation reported in the present work closes this loophole completely because the user does not report any losses to the bank. This is possible while still guaranteeing the protocol to work correctly given the high efficiency ($88.24\%$) of our setup. We note that Schiansky et al. [35] did not discuss the problem of multiphoton attacks in their implementation, nor mention any countermeasures employed. Without such countermeasures, user privacy is compromised.
@@
Our scheme assumes secure and authenticated classical channels among agents of the same party. These are straightforward to implement with pre-distributed secret keys. Our scheme can also be straightforwardly adapted to prevent any third party from impersonating the user or the bank in any of the communications between the bank and the user (using previously distributed secret keys, for instance). It can also be straightforwardly adapted to prevent any third party from impersonating the bank in the quantum communication, which occurs only once and is from the bank to the user. For example, after transmitting the quantum states to the user in the quantum token generation stage, the bank can indicate the states prepared, the bases used for preparation, and the labels for a random subset, using a secure and authenticated classical channel (implemented with previously distributed secret keys, for instance), which effectively allows the user to authenticate that the quantum communication was from the bank. The quantum channel does not need to be assumed secure, because the security properties of user privacy and unforgeability hold straightforwardly as claimed even if some third party is interfering in the quantum channel. Since our scheme satisfies user privacy against a dishonest bank, it also trivially satisfies user privacy against a third party trying to impersonate the bank. Since our scheme satisfies unforgeability against a dishonest user, it also trivially satisfies unforgeability against a third party trying to impersonate the user.
@@ 6 Security proof
which is an upper bound on the probability that a quantum state that Bob sends Alice has a dimension greater than two (corresponding to more than one qubit) or that it deviates from the intended BB84 state by an angle greater than $\theta$ on the Bloch sphere. In our updated unforgeability proof we assume, very conservatively, that an arbitrarily powerful dishonest Alice can perfectly read the bits that Bob encodes in quantum states with dimension greater than two (in multiple qubit states) and in qubit states that deviate from BB84 states by an angle greater than $\theta$. In practice the information she can extract from these infrequent “rogue” pulses will generally be far more limited.
@@ 6.3 Proof of unforgeability (proof of theorem 1)
6.4 No losses are reported: all received quantum states are qubits
@@
where $\rho^{k}=\sum_{i=1}^{4}q^{k}_{i}\rho^{k}_{i}$, $r^{k}_{i}=\frac{q^{k}_{i}+q^{k}_{i+1}}{2}$, and $\chi^{k}_{i}=\frac{q^{k}_{i}\rho^{k}_{i}+q^{k}_{i+1}\rho^{k}_{i+1}}{q^{k}_{i}+q^{k}_{i+1}}$, for all $i=1,2,3,4$. We note that $\{r^{k}_{i}\}_{i=1}^{4}$ is a probability distribution and that $\chi^{k}_{i}$ is a density matrix, for all $i=1,2,3,4$. We also note that $\sum_{i=1}^{4}r^{k}_{i}\chi^{k}_{i}=\rho^{k}$. Thus, we see that $P^{k,\text{win}}_{yz}$ is twice the probability $P^{k,\text{discr}}_{yz}$ to succeed in a quantum state discrimination task given by the ensemble $\{r^{k}_{i},\chi^{k}_{i}\}_{i=1}^{4}$ conditioned on the extra state being $\phi_{z}$ and the extra outcome being $y$:
@@
6.5 No losses are reported: a fraction of quantum states are multiple qubits
@@
Therefore, from (170), (6.5) and (6.5), we have
@@
6.6 No losses are reported: allowing $\theta_{k}>\theta$ with a non-zero probability $P_{\theta}$
@@
6.7 The general case: losses are reported
@@
We note that the bounds (6.7) and (6.7) hold for any values of the bit $c$, the set $\Lambda$ satisfying (181) and any other extra variables included in $v$. Thus, from (182), (6.7) and (6.7), we obtain that Alice’s cheating probability satisfies the bound
@@
6.8 Proof for the quantum token scheme $\mathcal{QT}_{1}$
@@ 7 Extension to a global network
We showed our implementation to be $\epsilon_{\text{unf}}-$unforgeable with $\epsilon_{\text{unf}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}5.52\times 10^{-9}}$. A very weak bound on the unforgeability of $\mathcal{QT}_{2}^{M}$ can be obtained [14] by noting that, to successfully forge, Alice needs to present a validated token at least one of the $\frac{1}{2}2^{M}(2^{M}-1)$ pairs of distinct presentation points, which means her forging probability is bounded by

### v2

@@ 5 Related work and extensions of our schemes
If implemented in a scenario in which transaction times and positions are critical, users/clients of Ref. [38] would need to rely on clock synchronization and position authentication, using networks which are potentially insecure. As noted above, these issues potentially affect the correctness of any quantum token scheme, including the present implementation of ours and that of Ref. [38], when position and time are relevant. If the client’s coordinates are incorrect, she may present a token in an incorrect region, i.e. one where no merchant or bank agent is present, or in a valid but unintended region. This means she may spend her token to acquire resources in the wrong space-time region, which may be to her disadvantage (for example, if she has a time- and location-dependent trading strategy). It also affects the privacy of any quantum token scheme when position and time are relevant. The client may present quantum money tokens in a region in the causal past of her intended region, which would allow the spoofer (e.g. the bank, or a third party) to learn her intended presentation region. Alternatively, she may present tokens in a region space-like separated from her intended region, which allows the spoofer exploitable information about her trading strategy in regions where she requires such information to be unavailable. Again, this may be disadvantageous when her trading strategy and those of adversaries (such as the spoofer) are time- and location-dependent and utilize all locally available information. We emphasize again that secure position authentication and time synchronization are solvable problems – assuming that some data, some locations and a reference clock can be trusted – although they are not experimentally addressed in our present implementation, nor in Ref. [38], nor indeed, as far as we are aware, in any implementation to date in the field of relativistic quantum cryptography (see e.g. [31, 32, 52, 53, 54]). Of course, in scenarios in which transaction times and positions are not important, users of our scheme can label possible recipients of token presentations in ways that are not related to any purported space time coordinates. In this case they need not rely on secure clock synchronisation or position authentication.
@@
Neither the present scheme nor that of Ref. [38] require space-like separated presentation regions or other space-time constraints to ensure unforgeability. The reason we impose these constraints in the present implementation is to guarantee that the implementation demonstrates an advantage over classical alternatives.
@@
Our scheme assumes secure and authenticated classical channels among agents of the same party. These can be straightforwardly and efficiently implemented with pre-distributed secret keys. Our scheme can also be straightforwardly adapted to prevent any third party from impersonating the user or the bank in any of the communications between the bank and the user (using previously distributed secret keys, for instance). It can also be straightforwardly adapted to prevent any third party from impersonating the bank in the quantum communication, which occurs only once and is from the bank to the user. For example, after transmitting the quantum states to the user in the quantum token generation stage, the bank can indicate the states prepared, the bases used for preparation, and the labels for a random subset, using a secure and authenticated classical channel (implemented with previously distributed secret keys, for instance), which effectively allows the user to authenticate that the quantum communication was from the bank. The quantum channel does not need to be assumed secure, because the security properties of user privacy and unforgeability hold straightforwardly as claimed even if some third party is interfering in the quantum channel. Since our scheme satisfies user privacy against a dishonest bank, it also trivially satisfies user privacy against a third party trying to impersonate the bank. Since our scheme satisfies unforgeability against a dishonest user, it also trivially satisfies unforgeability against a third party trying to impersonate the user.
@@ 6 Security proof
which is an upper bound on the probability that a quantum state that Bob sends Alice has dimension greater than two (corresponding to more than one qubit) or that it deviates from the intended BB84 state by an angle greater than $\theta$ on the Bloch sphere. In our updated unforgeability proof we assume, very conservatively, that an arbitrarily powerful dishonest Alice can perfectly read the bits that Bob encodes in quantum states with dimension greater than two (in multiple qubit states) and in qubit states that deviate from BB84 states by an angle greater than $\theta$. In practice the information she can extract from these inrequent “rogue” pulses will generally be far more limited.
@@ 6.3 Proof of unforgeability (proof of theorem 1)
6.3.1 No losses are reported: all received quantum states are qubits
@@
where $\rho^{k}=\sum_{i=1}^{4}q^{k}_{i}\rho^{k}_{i}$, $r^{k}_{i}=\frac{q^{k}_{i}+q^{k}_{i+1}}{2}$, and $\chi^{k}_{i}=\frac{q^{k}_{i}\rho^{k}_{i}+q^{k}_{i+1}\rho^{k}_{i+1}}{q^{k}_{i}+q^{k}_{i+1}}$, for all $i=1,2,3,4$. We note that $\{r^{k}_{i}\}_{i=1}^{4}$ is a probability distribution and that $\chi^{k}_{i}$ is a density matrix, for all $i=1,2,3,4$. We also note that $\sum_{i=1}^{4}r^{k}_{i}\chi^{k}_{i}=\rho^{k}$. Thus, we see that $P^{k,\text{win}}_{yz}$ is twice the probability $P^{k,\text{discr}}_{yz}$ to succeed in a quantum state discrimination task given by the ensemble $\{r^{k}_{i},\chi^{k}_{i}\}_{i=1}^{4}$ conditioned on on the extra state being $\phi_{z}$ and the extra outcome being $y$:
@@
6.3.2 No losses are reported: a fraction of quantum states are multiple qubits
@@
Therefore, from (170), (6.3.2) and (6.3.2), we have
@@
6.3.3 No losses are reported: allowing $\theta_{k}>\theta$ with a non-zero probability $P_{\theta}$
@@
6.3.4 The general case: losses are reported
@@
We note that the bounds (6.3.4) and (6.3.4) hold for any values of the bit $c$, the set $\Lambda$ satisfying (181) and any other extra variables included in $v$. Thus, from (182), (6.3.4) and (6.3.4), we obtain that Alice’s cheating probability satisfies the bound
@@
6.3.5 Proof for the quantum token scheme $\mathcal{QT}_{1}$
@@ 7 Extension to a global network
We showed our implementation to be $\epsilon_{\text{unf}}-$unforgeable with $\epsilon_{\text{unf}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}5.52\times 10^{-9}}$. A very weak bound on the unforgeability of $\mathcal{QT}_{2}^{M}$ can be obtained [14] by noting that, to successfully forge, Alice needs to present a validated token at at least one of the $\frac{1}{2}2^{M}(2^{M}-1)$ pairs of distinct presentation points, which means her forging probability is bounded by
@@
8 Acknowledgements
We thank Mathieu Bozzio, Tobias Guggemos, Peter Schiansky and Philip Walther for very helpful discussions relating to Section 5 of this SI.